# No-prior Bayes reIMagined: probabilistic approximations of possibilistic inferential models

### v2

@@ 4.3 Marginalization risks, or lack thereof
Fraser’s “risks” primarily concern marginalization (e.g., Dawid et al., 1973; Fraser, 2011a ; Balch et al., 2019). Roughly, whatever desirable properties a no-prior Bayes solution for $\Theta$ has, the corresponding no-prior Bayes solution for $\Phi=m(\Theta)$ obtained by probabilistic marginalization generally doesn’t share those desirable properties. When Schweder and Hjort, (2013) warn that “joint [confidence distributions] should not be sought, we think, since they might easily lead the statistician astray,” they’re concerned about users being tempted to do these familiar probabilistic operations, thus creating a risk of unreliability. Fraser, 2011a and Fraser et al., (2016) highlight non-linearity in $m$ as a major contributor to this risk; see, also, Martin, 2023a ; Martin, 2024b and the discussion below.
There are different ways to transform a quantification of uncertainty about $\Theta$ to a one about $\Phi=m(\Theta)$. Start with a confidence set $C_{\alpha}(x)$. The image $C_{\alpha}^{m}(x)=m\{C_{\alpha}(x)\}$ is a confidence set for $\Phi$ as a result of the following implication:
@@
That is, $\mathsf{Q}_{x}^{m}$ is inclined to assign more of its mass to the marginal confidence sets than $\mathsf{Q}_{x}$ assigned to the original confidence sets. So, if $\mathsf{Q}_{x}$ was a confidence distribution (CD), i.e., it assigns probability $1-\alpha$ to $C_{\alpha}(x)$ for each $\alpha\in[0,1]$, then (12) suggests calling $\mathsf{Q}_{x}^{m}$ an over-confidence distribution (oCD). Illustrations of this over-confidence phenomenon are presented in Appendix D. This can be problematic because surely the user of $\mathsf{Q}_{x}^{m}$ will disregard the safe confidence set $C_{\alpha}^{m}(x)$ and, instead, construct his/her own quantification of uncertainty directly based on the over-confident $\mathsf{Q}_{x}^{m}$, risking unreliability. This begs the question: when does $\mathsf{Q}_{x}^{m}$’s credible set agree with the safe $C_{\alpha}^{m}(x)$?
Let $\overline{\Pi}_{x}$ be the possibilistic IM as described above, quantifying uncertainty about $\Theta$, given $X=x$. One way—but not the only way, see Section 6—to carry out possibilistic marginalization is by applying the general extension operation (e.g., Zadeh, 1975a ; Zadeh, 1975b ; Zadeh, 1978). Extension yields a new marginal IM $\overline{\Pi}_{x}^{m}$ for $\Phi=m(\Theta)$ with possibility contour
@@
That the possibilistic marginalization strategy appears organically in the common-sense treatment of confidence sets is no coincidence; see Martin, 2021a . The inner probabilistic approximations of $\overline{\Pi}_{x}$ are confidence distributions, and these are characterized as mixtures of kernels supported on the boundaries of the confidence sets. Then the derived marginal distribution $\mathsf{Q}_{x}^{m}$ of $\Phi=m(\Theta)$ clearly also can be characterized by a similar mixture, but $\mathsf{Q}_{x}^{m}$ would only be the inner probabilistic approximation of $\overline{\Pi}_{x}^{m}$ if the boundaries $\partial C_{\alpha}(x)$ of the original confidence sets get mapped via $m$ to the boundaries $\partial C_{\alpha}^{m}(x)$ of the marginal confidence sets. Define $m$ to be boundary-preserving if $m\{\partial C_{\alpha}(x)\}=\partial C_{\alpha}^{m}(x)$ for each $\alpha\in[0,1]$. Among other forms of mappings, linear transformations are boundary preserving because they attain their extrema on the boundary of their domain.
@@
If $\mathsf{Q}_{x}$ is an inner probabilistic approximation of $\overline{\Pi}_{x}$, and $m$ is a boundary-preserving mapping, then the marginal distribution $\mathsf{Q}_{x}^{m}$ for $\Phi=m(\Theta)$ derived from $\mathsf{Q}_{x}$ is the inner probabilistic approximation of $\overline{\Pi}_{x}^{m}$.
That the probability calculus leads to reliable marginalization effectively only when the transformation is linear doesn’t imply that reliable marginalization is out of reach in the non-linear case. It just means non-probabilistic marginalization strategies are needed. This isn’t unfamiliar: Bayesians choose default prior distributions that are tailored to the particular inferential target $\Phi$ (e.g., Tibshirani, 1989; Datta and Ghosh, 1995; Bernardo, 1979; Berger et al., 2009), and confidence and fiducial distribution users carry out various manipulations akin to those used to construct approximate pivotal quantities. The above insights motivate the following suggestion: first carry out the appropriate possibilistic marginalization to get $\pi_{x}^{m}$ and $\overline{\Pi}_{x}^{m}$, and then construct the inner probabilistic approximation of this. An example of this using the extension-based marginalization is shown below. Extension-based marginalization is a simple and very general, but is not the most efficient. Typically, sharper inferences can be drawn when marginalization is carried out by “validifying” a relative profile likelihood. Details of this profile-based approach will be presented in the context of a concrete example in Section 6 below.
@@
Figure 5: Marginal inference results for $\Phi=\cos(a\Theta)$ with $a=1$ in the directional data example. Panel (a) shows the extension-based marginal IM contour for $\Phi$; Panel (b) shows the Bayesian marginal posterior distribution for $\Phi$; and Panel (c) shows the non-credibility function for the marginal Bayesian posterior (dashed) and the marginal IM’s inner probabilistic approximation (solid).
Next, take $a=1.5$. Then the aforementioned “effectively boundary-preserving” property no longer holds and, surprisingly, this seemingly small change in the quantity of interest has non-negligible impact on the results. Figure 6 shows the same plots as in Figure 5, this time for $a=1.5$. The key observation is that, while the general shape of the marginal IM contour and the Bayes posterior distribution were similar in Figure 5, they are quite different in Figure 6; this is a consequence of the entirely different style of marginalization, from $\Theta$ to $\Phi$, with the IM’s possibilistic marginalization preserving the validity properties enjoyed by both approaches’ inference on $\Theta$ whereas the Bayesians’ probabilistic marginalization doesn’t. The IM’s inner probabilistic approximation (not shown) looks similar to the Bayesian posterior, but it’s not the same. Indeed, Figure 6(c) shows the non-credibility for the Bayes posterior distribution and the IM’s inner probabilistic approximation, and there are clear signs of the Bayesians’ over-confidence. The IM’s inner approximation also shows (milder) signs of over-confidence, but there’s a perfectly reasonable explanation for this. Recall from Section 3 that exact inner probabilistic approximations need not exist, e.g., if the possibility contour is bounded away from 0. In Figure 6(a), note that the contour is in fact bounded away from 0 and, consequently, the sets $C_{\alpha}^{m}(x)$ are equal to the full space $[-1,1]$ for sufficiently small $\alpha$. Since any probability distribution $\mathsf{Q}_{x}^{m}$ supported on $[-1,1]$ must assign probability $1>1-\alpha$ to $C_{\alpha}^{m}(x)=[-1,1]$ for all sufficiently small $\alpha$, there is no $\mathsf{Q}_{x}^{\star}$ that satisfies (9) for all $\alpha$. But the IM’s pseudo inner probabilistic approximation—“pseudo” because the inner probabilistic approximation strictly doesn’t exist—quickly corrects for this over-confidence at small $\alpha$ and achieves the objective (9) for all $\alpha$ that are not too small.
@@
Figure 6: Marginal inference results for $\Phi=\cos(a\Theta)$ with $a=1.5$ in the directional data example. Panel (a) shows the extension-based marginal IM contour for $\Phi$; Panel (b) shows the Bayesian marginal posterior distribution for $\Phi$; and Panel (c) shows the non-credibility function for the marginal Bayesian posterior (dashed) and the marginal IM’s inner probabilistic approximation (solid).
@@ 5 Computation
Until recently, only naive and relatively inefficient strategies for computing the IM contour were available. In particular, the go-to strategy was to approximate $\pi_{x}$ via
$\pi_{x}(\theta)\approx\frac{1}{M}\sum_{m=1}^{M}1\{R(X_{m,\theta},\theta)\leq R(x,\theta)\},\quad\theta\in\mathbb{T},$ (14)
@@
where $X_{m,\theta}$ are independent copies of the data $X$, drawn from $\mathsf{P}_{\theta}$, for $m=1,\ldots,M$. The above computation is feasible at a few different $\theta$ values, but all the practically relevant calculations—e.g., identifying the confidence set in (5)—require evaluation over a sufficiently fine grid covering the relevant portions of $\mathbb{T}$, and this is expensive. Recently, Martin, (2025) developed a new and more efficient strategy for doing the relevant IM computations that replaces most of the naive contour evaluations as in (14) with Monte Carlo sampling from a sort of “posterior distribution.” That distribution is, in fact, the inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ of $\overline{\Pi}_{x}$ defined in Section 3 above. Therefore, the same computational strategy for (approximately) sampling from $\mathsf{Q}_{x}^{\star}$ advanced in Martin, (2025) can be reIMagined here as an implementation of the new no-prior Bayes solution.
@@
The strategy in Martin, (2025) is, roughly, to stitch together a suitable collection of Gaussian possibilistic IMs. While the “stitched” collection of Gaussian possibilities isn’t a Gaussian possibility, the use of a basic Gaussian building block can be justified by the asymptotic normality result in Section 4.2. Interest here is solely in the inner probabilistic approximation, so it’s not necessary for me to describe all the details involved in the “stitched IM.” For $\xi$ a generic $D$-vector of positive numbers, consider a collection of Gaussian confidence/level sets
$C_{\alpha}^{\xi}(x)=\{\theta:(\theta-\hat{\theta}_{x})^{\top}\,J_{x}^{\xi}\,(\theta-\hat{\theta})\leq\chi_{D,1-\alpha}^{2}\},$
@@
where $\chi_{D,1-\alpha}^{2}$ is the $1-\alpha$ quantile of the ${\sf ChiSq}(D)$ distribution, and $J_{x}^{\xi}$ is an embellished version of the observed Fisher information matrix with spread at least partially controlled by the parameters in $\xi$. Specifically, if $J_{x}$ has spectral decomposition $J_{x}=E\Lambda E^{\top}$, then the embellished version $J_{x}^{\xi}$ is defined as
$J_{x}^{\xi}=E\,\text{diag}(\xi^{-1})\,\Lambda\,\text{diag}(\xi^{-1})\,E^{\top},$ (15)
@@
where $\text{diag}(\xi^{-1})$ is the diagonal matrix with elementwise reciprocal of the $\xi$ vector on the diagonal. Then $\xi$ controls confidence set’s size by manipulating the eigenvalues of the underlying observed Fisher information matrix, e.g., a large (resp. small) entry in $\xi$ stretches (resp. contracts) the level set in the direction of the corresponding eigenvector. Theorem 4 says that the original $C_{\alpha}(x)$ will be roughly elliptical, so it’s safe to assume that there exists a vector $\xi$ such that the following inclusion holds snugly:
$C_{\alpha}(x)\subseteq C_{\alpha}^{\xi}(x).$ (16)
@@
Of course, this vector would generally depend on $(x,\alpha)$, so I’ll denote this as $\xi(x,\alpha)$. A cartoon depiction of the proposed strategy is shown in Figure 7. The motivation for bounding $C_{\alpha}(x)$ inside $C_{\alpha}^{\xi}(x)$ is that it’s not possible to describe the former exactly but the latter is very simple, at least for fixed $\xi$. The data-driven algorithm recommended in Martin, (2025), relying on the foundations laid in Cella and Martin, (2024), for selecting $\xi(x,\alpha)$ is presented in Appendix E. Once this $\xi(x,\alpha)$ has been determined, the inner probabilistic approximation uses the kernel
$\mathsf{K}_{x}^{\alpha}={\sf Unif}\{\partial C_{\alpha}^{\xi(x,\alpha)}(x)\},\quad\alpha\in(0,1).$
@@
Thanks to the direction of the inclusion in (16), the proposed implementation of the inner probabilistic approximation is conservative. So, the relevant properties enjoyed in theory by the inner probabilistic approximation of the IM should also be at least approximately enjoyed by the practical proposal above; see Section 6.
[…]
@@
Figure 7: Approximating the IM contour’s $\alpha$-level set $C_{\alpha}(x)$ (gray) by an ellipse $C_{\alpha}^{\xi}(x)$ with a “good” choice of $\xi$ (solid) and with a “bad” choice (dashed).

### v3

@@ 4.3 Marginalization risks, or lack thereof
Fraser’s “risks” primarily concern marginalization (e.g., Dawid et al., 1973; Fraser, 2011a ; Balch et al., 2019). Roughly, whatever desirable properties a no-prior Bayes solution for $\Theta$ has, the corresponding no-prior Bayes solution for $\Phi=m(\Theta)$ obtained by probabilistic marginalization generally doesn’t share those desirable properties. When Schweder and Hjort, (2013) warn that “joint [confidence distributions] should not be sought, we think, since they might easily lead the statistician astray,” they’re concerned about users being tempted to do these familiar probabilistic operations, thus creating a risk of unreliability. Fraser, 2011a and Fraser et al., (2016) highlight non-linearity in $m$ as a major contributor to this risk; see, also, Martin, 2024b and the discussion below.
It’s important to emphasize that no probabilistic uncertainty quantification is safe from all such risks, including my proposed inner probabilistic approximations. So, the goal of the present section is simply to understand how marginalization with these approximations works and to identify which marginal inferences are safe. To be safe from all such risks, one must break from the familiar probabilistic uncertainty quantification in one way or another: Grünwald, (2018) suggests explicitly restricting one’s probabilistic inferences to those that are safe, whereas I’ve suggested relaxing from probabilities to possibilities so that all inferences are safe (Martin, 2025b ).
There are different ways to transform a quantification of uncertainty about $\Theta$ to a one about $\Phi=m(\Theta)$. Start with a confidence set $C_{\alpha}(x)$. The image $C_{\alpha}^{m}(x)=m\{C_{\alpha}(x)\}$ is a confidence set for $\Phi$ because
@@
That is, $\mathsf{Q}_{x}^{m}$ is inclined to assign more of its mass to the marginal confidence sets than $\mathsf{Q}_{x}$ assigned to the original confidence sets. So, if $\mathsf{Q}_{x}$ was a confidence distribution (CD), i.e., it assigns probability $1-\alpha$ to $C_{\alpha}(x)$ for each $\alpha\in[0,1]$, then (12) suggests calling $\mathsf{Q}_{x}^{m}$ an over-confidence distribution (oCD). Illustrations of this over-confidence phenomenon are presented in Appendix E of the supplement. This can be problematic because surely the user of $\mathsf{Q}_{x}^{m}$ will disregard the safe confidence set $C_{\alpha}^{m}(x)$ and, instead, construct his/her own quantification of uncertainty directly based on the over-confident $\mathsf{Q}_{x}^{m}$, risking unreliability. This begs the question: when does $\mathsf{Q}_{x}^{m}$’s credible set agree with the safe $C_{\alpha}^{m}(x)$?
Let $\mathsf{\Pi}_{x}$ be the possibilistic IM as described above, quantifying uncertainty about $\Theta$, given $X=x$. One way—but not the only way, see Section 6—to carry out possibilistic marginalization is by applying the general extension operation (e.g., Zadeh, 1975, 1978). Extension yields a new marginal IM $\mathsf{\Pi}_{x}^{m}$ for $\Phi=m(\Theta)$ with possibility contour
@@
That the possibilistic marginalization strategy appears organically in the common-sense treatment of confidence sets is no coincidence; see Martin, 2021a . The inner probabilistic approximations of $\mathsf{\Pi}_{x}$ are confidence distributions, and these are characterized as mixtures of kernels supported on the boundaries of the confidence sets. Then the derived marginal distribution $\mathsf{Q}_{x}^{m}$ of $\Phi=m(\Theta)$ clearly also can be characterized by a similar mixture, but $\mathsf{Q}_{x}^{m}$ would only be the inner probabilistic approximation of $\mathsf{\Pi}_{x}^{m}$ if the boundaries $\partial C_{\alpha}(x)$ of the original confidence sets get mapped via $m$ to the boundaries $\partial C_{\alpha}^{m}(x)$ of the marginal confidence sets. Define $m$ to be boundary-preserving if $m\{\partial C_{\alpha}(x)\}=\partial C_{\alpha}^{m}(x)$ for each $\alpha\in[0,1]$. Among other forms of mappings, linear transformations are boundary preserving because they attain their extrema on the boundary of their domain.
@@
If $\mathsf{Q}_{x}$ is an inner probabilistic approximation of $\mathsf{\Pi}_{x}$, and $m$ is a boundary-preserving mapping, then the marginal distribution $\mathsf{Q}_{x}^{m}$ for $\Phi=m(\Theta)$ derived from $\mathsf{Q}_{x}$ is an inner probabilistic approximation of $\mathsf{\Pi}_{x}^{m}$.
That the probability calculus leads to reliable marginalization effectively only when the transformation is linear doesn’t imply that reliable marginalization is out of reach in the non-linear case. It just means non-probabilistic marginalization strategies are needed. This isn’t unfamiliar: Bayesians choose default prior distributions tailored to the particular target $\Phi$ (e.g., Bernardo, 1979; Datta and Ghosh, 1995), and fiducialists apply various manipulations akin to those used to construct approximate pivotal quantities. These insights motivate the following suggestion: first carry out the appropriate possibilistic marginalization to get $\pi_{x}^{m}$ and $\mathsf{\Pi}_{x}^{m}$, and then construct the inner probabilistic approximation. An example of this using the extension-based marginalization is shown below. Extension-based marginalization is a simple and very general approach, but is not the most efficient. Typically, sharper inferences can be drawn when marginalization is carried out by “validifying” a relative profile likelihood. Details of this profile-based approach will be presented in the context of a concrete example in Section 6 below.
@@
Figure 5: Marginal inference results for $\Phi=\cos(a\Theta)$, with $a=1$, in the directional data example. Panel (a) shows the extension-based marginal IM contour for $\Phi$; Panel (b) shows the Bayesian marginal posterior distribution for $\Phi$; and Panel (c) shows the non-credibility function for the marginal Bayesian posterior (dashed) and the marginal IM’s inner probabilistic approximation (solid).
Next, take $a=1.5$. Then the “effectively boundary-preserving” property no longer holds and, surprisingly, this small change in the quantity of interest has non-negligible impact on the results. Figure 6 shows the same plots as in Figure 5, this time for $a=1.5$. The key observation is that, while the general shape of the marginal IM contour and the Bayes posterior distribution were similar in Figure 5, they are quite different in Figure 6; this is a consequence of the entirely different style of marginalization, from $\Theta$ to $\Phi$, with the IM’s possibilistic marginalization preserving the validity properties enjoyed by both approaches’ inference on $\Theta$ whereas the Bayesians’ probabilistic marginalization doesn’t. The IM’s inner probabilistic approximation (not shown) looks similar to the Bayesian posterior, but it’s not the same. Indeed, Figure 6(c) shows the non-credibility for the Bayes posterior distribution and the IM’s inner probabilistic approximation, and there are clear signs of the Bayesians’ over-confidence. The IM’s inner approximation also shows (milder) signs of over-confidence, but there’s a perfectly reasonable explanation for this. Recall from Section 3 that exact inner probabilistic approximations need not exist, e.g., if the possibility contour is bounded away from 0. In Figure 6(a), note that the contour is in fact bounded away from 0 and, consequently, the sets $C_{\alpha}^{m}(x)$ are equal to the full space [1,1][-1,1] for sufficiently small $\alpha$. Since any probability distribution $\mathsf{Q}_{x}^{m}$ supported on [1,1][-1,1] must assign probability $1>1-\alpha$ to $C_{\alpha}^{m}(x)=[-1,1]$ for all sufficiently small $\alpha$, there is no $\mathsf{Q}_{x}^{\star}$ that satisfies (9) for all $\alpha$. But the IM’s pseudo inner probabilistic approximation—“pseudo” because the inner probabilistic approximation strictly doesn’t exist—quickly corrects for this over-confidence at small $\alpha$ and achieves the objective (9) for all $\alpha$ that are not too small.
@@
Figure 6: Marginal inference results for $\Phi=\cos(a\Theta)$, with $a=1.5$, in the directional data example. Panel (a) shows the extension-based marginal IM contour for $\Phi$; Panel (b) shows the Bayesian marginal posterior distribution for $\Phi$; and Panel (c) shows the non-credibility function for the marginal Bayesian posterior (dashed) and the marginal IM’s inner probabilistic approximation (solid).
@@ 5 Computation
Until recently, only naive/inefficient strategies for computing the IM contour were available. In particular, the go-to strategy was to approximate $\pi_{x}$ via
$\pi_{x}(\theta)\approx\frac{1}{M}\sum_{m=1}^{M}1\{R(X_{m,\theta},\theta)\leq R(x,\theta)\},$ (14)
@@
where $X_{m,\theta}$ are independent copies of the data $X$, drawn from $\mathsf{P}_{\theta}$, for $m=1,\ldots,M$. The above computation is feasible at a few different $\theta$ values, but all the practically relevant calculations—e.g., identifying the confidence set in (5)—require evaluation over a sufficiently fine grid covering the relevant portions of $\mathbb{T}$, and this is expensive. Recently, Martin, 2025a developed a new and more efficient strategy for doing the relevant IM computations that replaces most of the naive contour evaluations as in (14) with Monte Carlo sampling from a sort of “posterior distribution.” That distribution is, in fact, the inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ of $\mathsf{\Pi}_{x}$ defined in Section 3 above. Therefore, the same computational strategy for (approximately) sampling from $\mathsf{Q}_{x}^{\star}$ advanced in Martin, 2025a can be reIMagined here as an implementation of the new no-prior Bayes solution.
@@
The strategy in Martin, 2025a is, roughly, to stitch together a suitable collection of Gaussian possibilistic IMs. While the “stitched” collection of Gaussian possibilities isn’t a Gaussian possibility, the use of a basic Gaussian building block can be justified by the asymptotic normality result in Section 4.2. Interest here is solely in the inner probabilistic approximation, so it’s not necessary for me to describe all the details involved in the “stitched IM.” For $\sigma$ a generic $D$-vector of positive numbers, consider a collection of Gaussian confidence/level sets
$C_{\alpha}^{\sigma}(x)=\{\theta:(\theta-\hat{\theta}_{x})^{\top}\,J_{x}^{\sigma}\,(\theta-\hat{\theta})\leq\chi_{D,1-\alpha}^{2}\},$
@@
where $\chi_{D,1-\alpha}^{2}$ is the $1-\alpha$ quantile of the ${\sf ChiSq}(D)$ distribution, and $J_{x}^{\sigma}$ is an embellished version of the observed Fisher information matrix with spread at least partially controlled by the parameters in $\sigma$. Specifically, if $J_{x}$ has spectral decomposition $J_{x}=E\Lambda E^{\top}$, then the embellished version $J_{x}^{\sigma}$ is defined as
$J_{x}^{\sigma}=E\,\text{diag}(\sigma^{-1})\,\Lambda\,\text{diag}(\sigma^{-1})\,E^{\top},$ (15)
@@
where $\text{diag}(\sigma^{-1})$ is the diagonal matrix with elementwise reciprocal of the $\sigma$ vector on the diagonal. Then $\sigma$ controls confidence set’s size by manipulating the eigenvalues of the underlying observed Fisher information matrix, e.g., a large (resp. small) entry in $\sigma$ stretches (resp. contracts) the level set in the direction of the corresponding eigenvector. Theorem 4 says that the original $C_{\alpha}(x)$ will be roughly elliptical, so it’s safe to assume that there exists a vector $\sigma$ such that the following inclusion holds snugly:
$C_{\alpha}(x)\subseteq C_{\alpha}^{\sigma}(x).$ (16)
@@
Of course, this vector $\sigma$ generally depends on $(x,\alpha)$, so I’ll denote this as $\sigma(x,\alpha)$. A cartoon depiction of the proposed strategy is shown in Figure 7. The motivation for bounding $C_{\alpha}(x)$ inside $C_{\alpha}^{\sigma}(x)$ is that it’s not possible to describe the former exactly but the latter is very simple, at least for fixed $\sigma$. The data-driven algorithm recommended in Martin, 2025a , relying on the foundations laid in Cella and Martin, (2025), for selecting $\sigma(x,\alpha)$ is presented in Appendix F in the supplement. Once this $\sigma(x,\alpha)$ has been determined, the inner probabilistic approximation uses the kernel
$\mathsf{K}_{x}^{\alpha}={\sf Unif}\{\partial C_{\alpha}^{\sigma(x,\alpha)}(x)\},\quad\alpha\in(0,1).$
@@
Thanks to the direction of the inclusion in (16), the proposed implementation of the inner probabilistic approximation is conservative. So, the relevant properties enjoyed in theory by the inner probabilistic approximation of the IM should be at least approximately enjoyed by the practical proposal above; see Section 6.
$\hat{\theta}_{x}$$C_{\alpha}(x)$$\partial C_{\alpha}^{\sigma(x,\alpha)}(x)$
@@
Figure 7: Approximating the IM contour’s $\alpha$-level set $C_{\alpha}(x)$ (gray) by an ellipse $C_{\alpha}^{\sigma}(x)$ with a “good” choice of $\sigma$ (solid) and with a “bad” choice of $\sigma$ (dashed).